Can the moment of force in a rotational motion be calculated by using both of these formulas: m.g.R(1-costheta) or m.g.R.theta


1
Expert's answer
2021-01-10T18:26:17-0500

Yes, for very low angles:


"1-\cos{\theta}=2\sin^2{(0.5\theta)}\\
\lim_{\theta\to0}{(1-\cos{\theta})}=0.5\theta"